# Advancing Universal Deep Learning for Electronic-Structure Hamiltonian Prediction of Materials

### v2

@@ 3.3 Training Loss Functions
The objective of training the neural network is to make the predicted Hamiltonian, denoted as $\widehat{\mathbf{H}}=\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$, approximate the ground truth $\Delta\mathbf{H}^{gt}$ as closely as possible. As illustrated in Fig. 2, to ensure that the predicted Hamiltonian can accurately derive down-stream physical quantities (such as band structures), we design a joint optimization strategy in both real space (R-space) and reciprocal space (k-space) for the neural network.
In R-space, the Hamiltonian and the corresponding trace quantity are jointly supervised. As outlined in Section 3.2, the trace quantity is used to supervise the non-linear SO(3)-invariant features, which contribute to constructing the non-linear SO(3)-equivariant features required for predicting the Hamiltonian. The R-space training loss function is defined as:
@@
where ${\mathbb{E}}_{\mathbf{R}}[\cdot]$ denotes the empirical expectation, $\lambda_{C}$, $\lambda_{R}$ are hyperparameters; $\mathbf{R}$ denotes the lattice vector connecting the reference unit cell and a neighboring unit cell; $\mathrm{loss}_{H}(\mathbf{R})$ and $\mathrm{loss}_{T}(\mathbf{R})$ denote the prediction losses of the Hamiltonian and the trace quantity in $\mathbf{R}$-space, respectively; and $\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})$ is a scaling factor designed to balance their relative contributions for stable training. The detailed forms of these terms are provided in Appendix A.3.
As analyzed in Appendix A.2, due to the error amplification mechanism associated with the ill-conditioned overlap matrix, even small numerical errors in $R$-space can be magnified in $k$-space, leading to deviations in downstream physical quantities. To mitigate this, we introduce $k$-space loss functions. Specifically, the spectrum is partitioned into a low-energy subspace $P$, which governs most physical properties, and a high-energy complement $Q$. While downstream phenomena are predominantly determined by $P$, an inaccurately predicted Hamiltonian may introduce spurious couplings between $P$ and $Q$. This can result in unphysical abrupt changes in band structures, which are referred to as “ghost states” (see Fig. 14 in Appendix A.8). Therefore, it is essential not only to emphasize accuracy in $P$ but also to maintain reasonable fidelity in $Q$ so that the erroneous $PQ$ couplings can be identified and suppressed. To this end, we incorporate differentiated weights for $P$ and $Q$ in the loss design, together with an explicit $PQ$ penalty that eliminates unphysical cross-subspace couplings and suppresses ghost states.
@@
where $\lambda_{P}$, $\lambda_{Q}$, and $\lambda_{PQ}$ are tunable hyperparameters that adjust the relative importance of the three loss terms, which respectively measure the errors in the $P$ subspace, the $Q$ subspace, and the combined $PQ$ joint subspace. The detailed formulations of these terms are provided in Appendix A.3.
@@ 4 Dataset
As broad-coverage open-source Hamiltonian datasets that use fine-grained orbital descriptions and include spin-orbit coupling (SOC) effects across a wide range of crystals are still rare, we construct one ourselves and contribute it to the community. Specifically, our dataset, called as Materials-HAM-SOC, contains 17,00017{,}000 material structures sampled from the Materials Project (Jain et al., 2013), with ground-truth Hamiltonians and band structures generated using the DFT software ABACUS (Li et al., 2016; Lin et al., 2023) and PYATB (Jin et al., 2023). It spans $68$ distinct elements from the first six rows of the periodic table and explicitly incorporates SOC effects. For these structures, a high-quality atomic orbital basis set (Lin et al., 2021), up to 4s2p2d1f orbitals for each element, is employed, providing a fine-grained representation of their electronic structure. The dataset contains all quantities required by our method, including atomic structures, zeroth-step Hamiltonians, self-consistent Hamiltonians, and overlap matrices. The dataset is partitioned into 12,000 structures for training, 2,000 for validation, and 3,000 for testing. For details of the dataset construction and comprehensive statistical summaries, please refer to Section A.4.
@@ 5 Empirical Study
We perform empirical studies on the Materials-HAM-SOC dataset. The implementation details of the network architecture and training configurations are provided in Appendix A.5.
@@
While mean absolute error (MAE) is a straightforward error metric, Hamiltonian prediction presents a unique gauge freedom: adding a global shift $\mu\mathbf{S}$, where $\mu$ is an arbitrary scalar and $\mathbf{S}$ is the overlap matrix, leaves all down-stream physical quantities unchanged (Wang et al., 2024b). This necessitates a gauge-invariant error metric for fair evaluation. To remove this gauge freedom, we adopt the Gauge MAE (Wang et al., 2024b) to our context:
@@
As shown in Table 1, the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ closely matches the self-consistent Hamiltonian $\mathbf{H}^{(T)}$ in the spin-flip submatrices ($\uparrow\downarrow$ and $\downarrow\uparrow$). Similarly, the imaginary parts of the spin-conserving submatrices ($\uparrow\uparrow$ and $\downarrow\downarrow$) also exhibit excellent agreement. In these components, the deviation between $\mathbf{H}^{(0)}$ and $\mathbf{H}^{(T)}$ is negligible, with errors reaching sub-$\mu eV$ level. Furthermore, in systems with time-reversal symmetry and real-valued atomic orbitals, which constitute the majority of practical cases, the real parts of the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are identical. This symmetry implies that the correction network only needs to predict the real part of the $\uparrow\uparrow$ block in $\Delta\mathbf{H}$, substantially reducing the number of matrix elements to be learned. The $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})$ for the real part of the $\uparrow\uparrow$ block is reduced by $96\%$ compared to $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$, yielding a much narrower numerical range for regression. This substantial reduction eases optimization by allowing the network to concentrate on physically meaningful residual corrections rather than reconstructing the entire Hamiltonian, thereby improving prediction accuracy across diverse atomic configurations. Finally, with the neural network correction applied, the errors for the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are substantially reduced, achieving an superior prediction accuracy: the overall Gauge MAE is 1.417meV1.417\ $\mathrm${meV}, closely matching the ground-truth labels obtained from DFT calculations.
In Fig. 3, we report a fine-grained evaluation of prediction accuracy by partitioning the test set into subsets defined by chemical elements. For each element, we gather all crystal structures that contain it and compute the mean error within this subset. The resulting per-element statistics are visualized on the periodic table, providing a clear view of how the model generalizes across chemically diverse systems. The analysis shows that for the most of elements the prediction errors are below 1.5meV1.5\ $\mathrm${meV}, confirming the robustness of our approach across a broad spectrum of the periodic table.
In material systems, the band structure serves as a fundamental representation of the electronic structure. We sample a variety of diverse examples from the testing set and compared the band structures derived from our predicted Hamiltonians with those obtained from DFT. The results show that the band structures predicted by our method closely match those obtained from DFT, highlighting its strong potential for practical applications in materials science. Details are provided in Appendix A.6. As detailed in Appendix A.7, we evaluate the end-to-end inference time of our method, including both the construction of the zeroth-step Hamiltonian and the neural network inference, and compare it with the DFT method. The results show that our approach delivers a substantial speedup.
As detailed in Appendix A.8, we conduct fine-grained ablation studies to assess the contributions of individual components in our framework. The results demonstrate that the physics-informed input descriptor $\mathbf{H}^{(0)}$, the correction-based regression target design, the TraceGrad mechanism, the ensemble strategy, and the joint $R$- and $k$-space training objective each provide significant error reductions, and that their combination corresponds to the best overall performance. In particular, incorporating the $k$-space loss markedly enhances the accuracy of the predicted band structures and effectively suppresses unphysical artifacts such as ghost states.
@@ 6 Conclusion
We advance universal Hamiltonian deep learning through both a new method and a new dataset . We propose NextHAM, a unified deep learning framework designed for accurate and generalizable prediction of electronic-structure Hamiltonians across the periodic table. First, we leverage zeroth-step Hamiltonians, constructed from initial charge densities, as both informative input features and output priors, allowing the model to perform correction-based regression and significantly reducing the learning complexity. Second, we present a Transformer-based neural architecture that enforces strict $E(3)$-equivariance while maintaining high expressive capacity, enabling accurate modeling of spatial symmetries in material systems. Third, we design a novel training objective that jointly optimizes the Hamiltonian prediction in both real space and reciprocal space, ensuring consistency with downstream physical quantities such as band structures. We also release Materials-HAM-SOC, a diverse-collection benchmark of 17,000 DFT-calculated material structures spanning six rows of the periodic table, with explicit spin–orbit coupling and high-resolution orbital representations, providing high-quality resources for training and evaluation. Empirically, NextHAM attains DFT-level accuracy for Hamiltonians and band structures while bringing substantial speedups over conventional DFT workflows, providing powerful tools to efficient simulation and design of new materials.
Appendix A Appendix
@@
A.1 Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
@@
Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has established itself as a foundational tool in modern electronic structure theory, with wide-ranging applications in condensed matter physics, quantum chemistry, and materials science. First developed in the 1960s by Hohenberg, Kohn, and Sham, DFT reformulates the many-electron problem by replacing the complex many-body wavefunction with the electron density $\rho(\mathbf{r})$ as the central variable. This shift dramatically simplifies the computational treatment of quantum systems while retaining the essential physics, making it feasible to study realistic systems under accepted computational cost. Over the years, DFT has become indispensable for tasks such as computing band structures and orbital energies, performing structural optimizations, and predicting a variety of electronic, magnetic, and optical properties. Its broad applicability and computational efficiency have cemented its role as a key methodology across multiple scientific domains.
@@
At the heart of density functional theory (DFT) lies the Kohn–Sham (KS) equation (Kohn & Sham, 1965), which reformulates the many-body electronic problem into a tractable set of single-particle equations:
@@
$\rho^{(0)}(\mathbf{r})\rightarrow V^{(0)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(0)}\rightarrow\psi^{(0)}_{n\mathbf{k}}(\mathbf{r})\rightarrow\rho^{(1)}(\mathbf{r})\rightarrow\cdots\rightarrow\rho^{(T)}(\mathbf{r})\rightarrow V^{(T)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(T)}$
@@
To address this challenge, recent approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023b; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) have adopted the deep graph learning paradigm to predict the self-consistent Hamiltonians. These methods bypass the iterative and computationally intensive matrix diagonalization steps in traditional DFT algorithms by directly predicting the final converged Hamiltonian matrix $H^{(T)}_{\alpha\beta}$ in a single forward pass. As shown in Eq. (11), the Hamiltonian matrix is inherently sparse: only pairs of atoms within a cutoff radius contribute non-zero elements. Therefore, the total number of Hamiltonian matrix elements that need to be computed scales with the number of local atomic pairs in the system, leading to a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ denotes the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis functions have finite spatial support, matrix elements vanish beyond a certain inter-atomic distance. In small systems where all atoms lie within each other’s cutoff radius, $\overline{E}\sim N$, and the total number of non-zero elements scales as $\mathcal{O}(N^{2})$. However, in sufficiently large systems, $\overline{E}$ saturates to a constant determined by local geometry, making the number of non-zero Hamiltonian elements scale linearly as $\mathcal{O}(N)$. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$ for large systems.

### v3

@@ 2.3 Training Loss Functions
The objective of training the neural network is to make the predicted Hamiltonian, denoted as $\widehat{\mathbf{H}}=\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$, approximate the ground truth $\Delta\mathbf{H}^{gt}$ as closely as possible. As illustrated in Fig. 2, to ensure that the predicted Hamiltonian can accurately derive down-stream physical quantities (such as band structures), we design a joint optimization strategy in both real space (R-space) and reciprocal space (k-space) for the neural network. Crucially, as detailed in Appendix G, our entire loss formulation is designed to rigorously resolve the gauge ambiguity Wang et al. (2024c), ensuring the uniqueness and physical consistency of the regression targets.
In R-space, the Hamiltonian and the corresponding trace quantity are jointly supervised. As outlined in Section 2.2, the trace quantity is used to supervise the non-linear O(3)-invariant features, which contribute to constructing the non-linear O(3)-equivariant features required for predicting the Hamiltonian. The R-space training loss function is defined as:
@@
where ${\mathbb{E}}_{\mathbf{R}}[\cdot]$ denotes the empirical expectation, $\lambda_{C}$, $\lambda_{R}$ are hyper-parameters; $\mathbf{R}$ denotes the lattice vector connecting the reference unit cell and a neighboring unit cell; $\mathrm{loss}_{H}(\mathbf{R})$ and $\mathrm{loss}_{T}(\mathbf{R})$ denote the prediction losses of the Hamiltonian and the trace quantity in $\mathbf{R}$-space, respectively; and $\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})$ is a scaling factor designed to balance their relative contributions for stable training. The detailed forms of these terms are provided in Appendix G.
As analyzed in Appendix F, due to the error amplification mechanism associated with the ill-conditioned overlap matrix, even small numerical errors in $R$-space can be magnified in $k$-space, leading to deviations in downstream physical quantities. To mitigate this, we introduce $k$-space loss functions. Specifically, the spectrum is partitioned into a low-energy subspace $P$, which governs most physical properties, and a high-energy complement $Q$. While downstream phenomena are predominantly determined by $P$, an inaccurately predicted Hamiltonian may introduce spurious couplings between $P$ and $Q$. This can result in unphysical abrupt changes in band structures, which are referred to as “ghost states” (see Fig. 15 in Appendix L). Therefore, it is essential not only to emphasize accuracy in $P$ but also to maintain reasonable fidelity in $Q$ so that the erroneous $PQ$ couplings can be identified and suppressed. To this end, we incorporate differentiated weights for $P$ and $Q$ in the loss design, together with an explicit $PQ$ penalty that eliminates unphysical cross-subspace couplings and suppresses ghost states.
@@
where $\lambda_{P}$, $\lambda_{Q}$, and $\lambda_{PQ}$ are tunable hyper-parameters that adjust the relative importance of the three loss terms, which respectively measure the errors in the $P$ subspace, the $Q$ subspace, and the combined $PQ$ joint subspace. The detailed formulations of these terms are provided in Appendix G.
@@ 3 Dataset
As broad-coverage open-source Hamiltonian datasets that use fine-grained orbital descriptions and include spin-orbit coupling (SOC) effects across a wide range of crystals are still rare, we construct one ourselves and contribute it to the community. Specifically, our dataset, called Materials-HAM-SOC, contains 17,00017{,}000 material structures sampled from the Materials Project (Jain et al., 2013), with ground-truth Hamiltonians and band structures generated using the DFT software ABACUS (Li et al., 2016; Lin et al., 2023) and PYATB (Jin et al., 2023). It spans more than $60$ distinct elements from the first six rows of the periodic table and explicitly incorporates SOC effects. For these structures, a high-quality atomic orbital basis set (Lin et al., 2021), up to 4s2p2d1f orbitals for each element, is employed, providing a fine-grained representation of their electronic structure. The dataset contains all quantities required by our method, including atomic structures, zeroth-step Hamiltonians, self-consistent Hamiltonians, and overlap matrices. The dataset is partitioned into 12,000 structures for training, 2,000 for validation, and 3,000 for testing. For details of the dataset construction and comprehensive statistical summaries, please refer to Section H.
@@ 4 Empirical Study
4.1 Statistical Results
We perform empirical studies on the Materials-HAM-SOC dataset. The implementation details of the network architecture and training configurations are provided in Appendix I.
@@ 4.1 Statistical Results
While mean absolute error (MAE) is a straightforward error metric, Hamiltonian prediction presents a unique gauge freedom: adding a global shift $\mu\mathbf{S}$, where $\mu$ is an arbitrary scalar and $\mathbf{S}$ is the overlap matrix, leaves all down-stream physical quantities unchanged (Wang et al., 2024c). This necessitates a gauge-invariant error metric for fair evaluation. To remove this gauge freedom, we adopt the Gauge MAE (Wang et al., 2024c) to our context:
@@
As shown in Table 1, the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ closely matches the self-consistent Hamiltonian $\mathbf{H}^{(T)}$ in the spin-flip submatrices ($\uparrow\downarrow$ and $\downarrow\uparrow$). Similarly, the imaginary parts of the spin-conserving submatrices ($\uparrow\uparrow$ and $\downarrow\downarrow$) also exhibit excellent agreement. In these components, the deviation between $\mathbf{H}^{(0)}$ and $\mathbf{H}^{(T)}$ is negligible, with errors reaching sub-$\mu eV$ level. Achieving SOC predictions with an accuracy of $\mu eV$ level holds significant value, as small differences in SOC energy can greatly impact the electronic structure of materials (Jing et al., 2025).
Furthermore, the $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})$ for the real part of the $\uparrow\uparrow$ block is reduced by $96\%$ compared to $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$, yielding a much narrower numerical range for regression. This substantial reduction eases optimization by allowing the network to concentrate on physically meaningful residual corrections rather than reconstructing the entire Hamiltonian, thereby improving prediction accuracy across diverse atomic configurations. In systems with time-reversal symmetry and real-valued atomic orbitals, which constitute the majority of practical cases, the real parts of the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are identical. This symmetry implies that the correction network only needs to predict the real part of the $\uparrow\uparrow$ block in $\Delta\mathbf{H}$, substantially reducing the number of matrix elements to be learned. Finally, with the neural network correction applied, the errors for the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are substantially reduced, achieving a superior prediction accuracy: the overall Gauge MAE is 1.417meV1.417\ $\mathrm${meV}, closely matching the ground-truth labels obtained from DFT calculations.
@@
In Fig. 3, we report a fine-grained evaluation of prediction accuracy by partitioning the test set into subsets defined by chemical elements. For each element, we gather all crystal structures that contain it and compute the mean error within this subset. The resulting per-element statistics are visualized on the periodic table, providing a clear view of how the model generalizes across chemically diverse systems. The analysis shows that for most of the elements, the prediction errors are below 1.5meV1.5\ $\mathrm${meV}, confirming the robustness of our approach across a broad spectrum of the periodic table.
For a more detailed analysis of the contributions of different components in our framework, we conduct fine-grained ablation studies, which are detailed in Appendix L. These studies show that the physics-informed input descriptor $\mathbf{H}^{(0)}$, the correction-based regression target design, the TraceGrad mechanism, the ensemble strategy, and the joint $R$- and $k$-space training objective each provide significant reductions in errors. The combination of all these components leads to the best overall performance, with notable improvements in both band structure prediction and the suppression of unphysical artifacts such as ghost states. We also compare our method with DeepH-E3 (Gong et al., 2023) and the original work of TraceGrad (Yin et al., 2025), demonstrating the significant superiority of our method. For details, please refer to Appendix M.
4.2 Case Study on Out-of-Distribution Generalization
As shown in Figure 7, our model is trained on a dataset that does not include structures containing the element Neon (Ne). However, in the testing set, a Ne-containing structure is included. Despite Ne being unseen during training, the model is able to generalize well to this new element, with testing error remaining very small, i.e., $0.1$ meV for the R-space MAE, as reported in Figure 3. This demonstrates the model’s ability to extrapolate knowledge learned from other elements to predict properties for unseen elements. This out-of-distribution generalization capability is a direct result of our approach using the zeroth-step Hamiltonian as a descriptor. Unlike traditional methods that rely on randomly initialized embeddings for each element, which cannot generalize to unseen elements not included in the training set, our model embeds physical information about the system’s electronic structure into a unified representation space using the zeroth-step Hamiltonian. This helps the model capture the relationships between different elements’ electronic structures, enabling it to generalize more effectively to out-of-distribution elements. This case study demonstrates the theoretical potential of our method’s generalizability to unseen elements. In future work, we will conduct further element-level out-of-distribution evaluations to more rigorously quantify this ability.
Figure 4: Panels a and b show the band-structure comparison among Ablation@Loss-k, the Full Method, and the Ground Truth DFT results. The setting Ablation@Loss-k drops the k-space loss from the Full Method. Panels c and d display the comparison of the real and imaginary parts of the optical conductivity along the x direction computed from the three Hamiltonians.
4.3 Case Study on $K$-space Loss Function
We present a representative case study regarding the effectiveness of the $k$-space loss function in the main text, with more cases available in Figure 15 of Appendix L. As shown in Fig. 4, in panel (a), the model is trained using only the $H(\mathbf{R})$-based loss defined in Eq. (1), ablating the $k$-space loss. In this case, the predicted band structure closely matches the DFT reference near the Fermi level for most $\mathbf{k}$-points. However, several isolated points exhibit abrupt discontinuities and deviate significantly from the ground-truth bands, which is a typical manifestation of ghost states. Notably, the MAE of the $H(\mathbf{R})$ matrix elements is only 0.53 meV. These sharp spectral anomalies show that even small MAE errors in real space ($H(\mathbf{R})$) can trigger ghost states in the resulting band structure.
This issue arises from the intrinsic numerical instability of the generalized eigenvalue problem, where the non-orthogonality of the orbital basis leads to pronounced error amplification. As we analyze in Appendix F, the sensitivity of both eigenvalues and eigenfunctions is enlarged by the factor $\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}$, implying that even small perturbations in the predicted $H(\mathbf{R})$ can cause significant deviations in the computed band energies, manifesting as ghost states in the spectrum.
In Fig. 4(b), we present the result from our Full Method, which includes the $k$-space loss. The predicted bands now align closely with the ground-truth results, and ghost states are nearly eliminated. The MAE of the real-space Hamiltonian $H(\mathbf{R})$ remains at $0.49\,\text{meV}$, similar to the case where only $\mathrm{loss}(\mathbf{R})$ was used, while the $k$-space loss $\mathrm{loss}(\mathbf{k})$ is reduced by more than $50\%$, substantially enhancing the overall quality of the predicted band structure.
Beyond band energies, the fidelity of the Hamiltonian is also reflected in wave-function-related physical observables. To further assess the influence of different loss functions on the wave-function accuracy, we compute the optical conductivity corresponding to the predictions in Figs. 4(a) and (b). The real and imaginary parts of the conductivity along the x direction, compared with the ground-truth results, are presented in Figs. 4(c) and (d). As demonstrated in these comparisons, the Hamiltonian trained with our Full Method yields substantially improved agreement with the reference conductivity. This result indicates that incorporating $\mathrm{loss}(k)$ not only suppresses ghost states under comparable $H(\mathbf{R})$ MAE, but also enhances the prediction of physical observables that depend sensitively on the wave-function quality.
@@ 5 Conclusion
We advance universal Hamiltonian deep learning through both a new method and a new dataset. We propose NextHAM, a unified deep learning framework designed for accurate and generalizable prediction of electronic-structure Hamiltonians across the periodic table. First, we leverage zeroth-step Hamiltonians constructed from initial charge densities as informative input features, facilitating the model to capture the intrinsic characteristics of electronic structures. Second, we present a Transformer-based neural architecture that enforces strict $E(3)$-equivariance while maintaining high expressive capacity, enabling accurate modeling of spatial symmetries in material systems. Third, we design a novel training objective that jointly optimizes the Hamiltonian prediction in both real space and reciprocal space, ensuring consistency with downstream physical quantities such as band structures. We also release Materials-HAM-SOC, a diverse-collection benchmark of 17,000 DFT-calculated material structures spanning six rows of the periodic table, with explicit spin–orbit coupling and high-resolution orbital representations, providing high-quality resources for training and evaluation. Empirically, NextHAM attains DFT-level accuracy for Hamiltonians and band structures while bringing substantial speedups over conventional DFT workflows, providing powerful tools to efficient simulation and design of new materials.
Acknowledgements
This work was supported by the Advanced Materials–National Science and Technology Major Project (Grant No. 2025ZD0618401), the National Natural Science Foundation of China (Grant No. 12134012), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500201), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301200). The numerical calculations were performed on the USTC High-Performance Computing facilities and Hefei advanced computing center.
Reproducibility statement
The codes and pre-trained models of NextHAM is available here: https://dzefile.hpccube.com:65011/efile/s/w/bmV4dGhhbQ==_0446ab078dd58ca3&, the extraction code is UUsg. The dataset Materials-SOC-HAM is available here: https://dzefile.hpccube.com:65011/efile/s/w/bmV4dGhhbQ==_c3c50c552df97ace&, the extraction code is DGEo. Details on loss functions are presented in Appendix G. Information regarding the dataset construction along with comprehensive statistical summaries can be found in Appendix H. The implementation details of the network architecture and the training setup are presented in Appendix I.
Ethics statement
This paper presents work whose goal is to advance the field of deep learning driven by materials research. Although our field has not yet exhibited direct negative social or ethical consequences, we recognize the importance of anticipating broader concerns. A key issue is the limited interpretability of deep learning systems, which obscures the reasoning behind their predictions and constrains their utility for gaining physical understanding. We stress the need for improving model interpretability, particularly regarding how physical knowledge is incorporated and represented, to guarantee both the accuracy of predictions and their applicability to a wide range of scientific problems.
@@ Acknowledgements
Appendix A Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
@@
Density Functional Theory (DFT) (Hohenberg and Kohn, 1964; Kohn and Sham, 1965) has established itself as a foundational tool in modern electronic structure theory, with wide-ranging applications in condensed matter physics, quantum chemistry, and materials science. First developed in the 1960s by Hohenberg, Kohn, and Sham, DFT reformulates the many-electron problem by replacing the complex many-body wavefunction with the electron density $\rho(\mathbf{r})$ as the central variable. This shift dramatically simplifies the computational treatment of quantum systems while retaining the essential physics, making it feasible to study realistic systems under accepted computational cost. Over the years, DFT has become indispensable for tasks such as computing band structures and orbital energies, performing structural optimizations, and predicting a variety of electronic, magnetic, and optical properties. Its broad applicability and computational efficiency have cemented its role as a key methodology across multiple scientific domains.
@@
At the heart of density functional theory (DFT) lies the Kohn–Sham (KS) equation (Kohn and Sham, 1965), which reformulates the many-body electronic problem into a tractable set of single-particle equations:
@@
$\rho^{(0)}(\mathbf{r})\rightarrow V^{(0)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(0)}\rightarrow\psi^{(0)}_{n\mathbf{k}}(\mathbf{r})\rightarrow\rho^{(1)}(\mathbf{r})\rightarrow\cdots\rightarrow\rho^{(T)}(\mathbf{r})\rightarrow V^{(T)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(T)}.$
@@
To address this challenge, recent approaches (Schütt et al., 2019; Unke et al., 2021; Li et al., 2022; Gong et al., 2023; Yu et al., 2023b; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) have adopted the deep graph learning paradigm to predict the self-consistent Hamiltonians. These methods bypass the iterative and computationally intensive matrix diagonalization steps in traditional DFT algorithms by directly predicting the final converged Hamiltonian matrix $H^{(T)}_{\alpha\beta}$ in a single forward pass. As shown in Eq. (11), the Hamiltonian matrix is inherently sparse: only pairs of atoms within a cutoff radius contribute non-zero elements. Therefore, the total number of Hamiltonian matrix elements that need to be computed scales with the number of local atomic pairs in the system, leading to a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ denotes the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis functions have finite spatial support, matrix elements vanish beyond a certain inter-atomic distance. In small systems where all atoms lie within each other’s cutoff radius, $\overline{E}\sim N$, and the total number of non-zero elements scales as $\mathcal{O}(N^{2})$. However, in sufficiently large systems, $\overline{E}$ saturates to a constant determined by local geometry, making the number of non-zero Hamiltonian elements scale linearly as $\mathcal{O}(N)$. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$ for large systems.